OpenStudy (anonymous):

The tennis team has won 10 matches and lost 5 matches. If the team continues to have the same ratio of wins to losses, how many wins will the team have after playing 45 matches?

OpenStudy (zarkon):

30

OpenStudy (anonymous):

10:5 = 2:1 total number of matches 10 + 5 + 45 = 60 so to keep the ratio the same you need 40 : 20 but the team has already won 10 matches so 40-10 = 30 so 30!

OpenStudy (zarkon):

or \[\frac{10}{15}=\frac{x}{45}\]

OpenStudy (zarkon):

\[x=45\frac{10}{15}=30\]